OpenStudy (anonymous):

How many hours is 2/3 of 1 hour and 36 minutes? PLEASE Help me answer it, i feel like im close...yet so far away!

OpenStudy (anonymous):

1 hour and 36 minutes is 96 minutes, agreed?

OpenStudy (anonymous):

corrrrrrect! and I get pretty stuck after that!

OpenStudy (anonymous):

okay. what is one-third of 96?

OpenStudy (anonymous):

okay when I type it in my calculator im getting 32! yea?

OpenStudy (anonymous):

ohh so like .5 hours?

OpenStudy (anonymous):

okay great! so one third of 96 is 32. what is two-thirds of 96?

OpenStudy (anonymous):

would i add 32 + 32? = 64?

OpenStudy (anonymous):

very good! so 2/3 of 96 minutes is 64 minutes. how many minutes in an hour?

OpenStudy (anonymous):

60 so its one hour 4 minutes?

OpenStudy (anonymous):

exactly! you can write your answer as 1 hour 4 minutes or (64/60) hours which is the same as 1.067 hours

OpenStudy (anonymous):

and I have another question I just need to figure out what method to use: Find the sum of the solutions of the equation: 2x\[^{2}\] +5x-3

OpenStudy (anonymous):

use factorisation

OpenStudy (anonymous):

2 * -3 = -6 think of combinations you can use to get 5x from factors of -6

OpenStudy (anonymous):

like -6 and 1?

OpenStudy (anonymous):

wait thats negative 5

OpenStudy (anonymous):

yup but -6 +1 = -5x, not 5x think of another combination

OpenStudy (anonymous):

okay are we adding the two factors? So like 2 +3= 5

OpenStudy (anonymous):

yes, 2+3 =5, but 2 times 3 is 6 we want -6 so 2 and 3 is not the correct combination

OpenStudy (anonymous):

okay, 6 plus -1 !! :)

OpenStudy (anonymous):

very good!

OpenStudy (anonymous):

go ahead and factorize it now

OpenStudy (anonymous):

omg that took a long time! haha okay i got (2x+6) and (x-1)

OpenStudy (anonymous):

no you should get (2x-1) and (x+3)

OpenStudy (anonymous):

then do i set them both to equal zero and solve or am i done with the solution

OpenStudy (anonymous):

the question asks to find the sum of the solution..

OpenStudy (anonymous):

\[2x ^{2}+5x-3 =0 \] \[2x ^{2}+6x-x-3 =0 \] \[2x (x+3)-1(x+3) =0\] \[(2x-1)(x-3)=0\] \[x = \frac{1}{2}, -3\]

OpenStudy (anonymous):

phenomenal, thank you thank you! and another thing when two slopes are parallel does that mean there slopes are the same?

OpenStudy (anonymous):

yes. it also means that there is no point of intersection.

OpenStudy (anonymous):

slopes cannot be parallel, only lines can be parallel. I assume you meant lines, not slopes in your question.

OpenStudy (anonymous):

oh okay okay yes I meant lines. for example 2/3 and 2/3 would be parallel lines correct?

OpenStudy (anonymous):

and if the question gives you a 4 quadrant grid and there is a triangle drawn on it and they ask how long is side AB , could you use the distance formula : "d= the square root of x2-x1) squared plus (y2-y1) squared ???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay thank you! You have truly been a help to me even so late! I am studying for my placement exam to College algebra at a community college and i think i got it i appreciate it!

OpenStudy (anonymous):

you are very knowledgeable, you should consider tutoring!

OpenStudy (anonymous):

lol, i wouldn't be any good at it. and you are welcome :)